# Bi-Directional Deep Contextual Video Compression

### v2

@@ II-A Deep Video Compression for P-Frame
For the second aspect, Li et al. [8] proposed a deep contextual video compression scheme, termed DCVC, which replaces the residual coding paradigm with a conditional coding paradigm. Instead of relying on the subtraction operation to reduce temporal redundancy, DCVC regards the temporal prediction as conditions so that the codec can learn how to make full use of the temporal prediction automatically. To further utilize temporal predictions, Sheng et al. [11] proposed DCVC-TCM, which not only uses temporal predictions at the entrance of the contextual encoder but also feeds the multi-scale temporal predictions into its intermediate locations. In addition, they also proposed to learn a temporal prior from the multi-scale temporal predictions to utilize temporal predictions in the entropy model. Based on DCVC-TCM, Li [9] proposed DCVC-HEM, which introduced a latent prior into the temporal entropy model to further utilize temporal correlation.
@@ IV-E Hierarchical Quality Structure
To achieve a hierarchical quality structure for B-frame coding, traditional video codecs commonly assign different QPs (quantization parameters) to video frames of different temporal layers [55]. Recent deep P-frame coding schemes [6, 10] periodically increase the quality of video frames by assigning a quality coefficient that varies periodically over the frame index to the Lagrangian multiplier $\lambda$ in the rate-distortion (R-D) loss function. Inspired by traditional video codecs and deep P-frame coding schemes, we propose to assign different quality coefficients $w_{t}$ for the B-frames of different temporal layers. As shown in Fig. 2, we set $w_{t}$ of the B-frames in Layer 1, 2, 3, 4, and 5 to [1.4, 1.4, 0.7, 0.5, 0.5], respectively. When calculating the loss functions to train our model, we multiply $w_{t}$ on the $\lambda$. In addition, to make our model adapt to different qualities better, a quality adaptor implemented by one $1\times 1$ convolutional layer is inserted before the feature extractor module of the temporal context mining. For B-frames of different temporal layers, different quality adaptors are selected. Note that the quality adaptors are included for training from the beginning.
We combine the hierarchical quality structure-based training strategy with the step-by-step training strategy. The detailed training strategies are listed in Table. I. For different numbers of training frames, the training reference structures are illustrated in Fig. 8. The definitions of the loss functions in different steps are described as follows.
@@
Different from existing deep P-frame coding schemes [11, 10, 12, 6], we do not calculate the average loss function of multiple frames. When using $L_{t}^{meD}$ or $L_{t}^{meRD}$ as the loss function, we only train the motion-related modules (Inter). When using $L_{t}^{recD}$ or $L_{t}^{recRD}$ as the loss function, we only train the context-related modules (Recon). When using $L_{t}^{all}$ as the loss function, we jointly train all the modules (All).
@@ V-A4 Test Configurations
For neural video coding models, we compare with existing state-of-the-art deep P-frame coding schemes, including DCVC-HEM [9], DCVC-DC [9], and DCVC-FM [12]. We do not compare with existing deep B-frame coding schemes since their best compression performance [14, 15] is still comparable with DCVC-HEM. The intra period and GOP size of our DCVC-B scheme and existing deep P-frame coding schemes are all set to 32. Different from [14] which compresses 97 frames of one video, we follow the existing deep P-frame coding schemes [10, 12] and compress 96 frames of one video to make a fair comparison. The structure of the first two GOPs is shown in Fig. 2. The structure of the last GOP is shown in Fig. 9, which is the same as the VTM-RA-GOP32 except that it has only 2 reference frames (VTM-RA-GOP32 has 4 reference frames). We use the same I-frame codec as DCVC-DC.
@@ V-B1 Comparison Results
We illustrate the RD-curves in terms of PSNR and MS-SSIM for different testing datasets in Fig. 10 and Fig. 11. We also report the corresponding BD-rate values in Table II and Table III. The anchor is HM-RA-GOP16. Negative values indicate bitrate saving compared with HM while positive values indicate bitrate increasing. The comparison results show that in terms of PSNR, our proposed DCVC-B scheme obtains an average –26.6% BD-rate reduction against the anchor HM-RA-GOP16, which even outperforms the SOTA deep P-frame coding schemes DCVC-DC (–24.4%) and DCVC-FM (–23.4%). For the HEVC Class D and Class E datasets, we even outperform VTM-RA-GOP32. Compared with the state-of-the-art deep B-frame coding schemes—B-CANF [14] and UCUV [15], which are only comparable to HM-RA-GOP16 and DCVC-HEM, our DCVC-B scheme achieves a large compression performance improvement. In terms of MS-SSIM, DCVC-B obtains an average –49.9% BD-rate reduction against HM-RA-GOP16 and even outperforms VTM-RA-GOP32 setting (–34.3%) over all the datasets. The subjective comparison results illustrated in Fig. 12 show that the reconstructed videos of our scheme can retain more details.
@@ V-C4 Effectiveness of Hierarchical Quality Structure
Figure 16: Frame quality and bitrate comparison between our proposed DCVC-B and DCVC-DC when the GOP size is set to 64.
@@
Datasets Sequences DCVC-DC DCVC-B Datasets Sequences DCVC-DC DCVC-B
HEVC Class B Cactus –31.0 –37.2 MCL-JCV videoSRC01 –17.4 –33.0
ParkScene –9.9 –30.8 videoSRC02 –16.0 0.4
Kimono1 –29.2 –29.7 videoSRC03 –23.4 –37.4
BQTerrace –12.2 –20.5 videoSRC04 –19.5 –13.4
BasketballDrive –26.3 –11.8 videoSRC05 –30.0 –11.3
HEVC Class C BQMall –20.3 –14.8 videoSRC06 –28.8 56.1
BasketballDrill –25.1 –26.9 videoSRC07 –29.9 –29.9
RaceHorses –22.0 –0.8 videoSRC08 –31.7 –20.8
PartyScene –15.4 –23.7 videoSRC09 –26.6 –40.1
HEVC Class D BasketballPass –36.4 –38.4 videoSRC10 -13.6 4.6
BQSquare –22.0 –42.4 videoSRC11 –35.5 –7.6
BlowingBubbles –24.5 –37.3 videoSRC12 –1.3 –31.3
RaceHorses –42.4 –34.9 videoSRC13 19.6 –18.0
HEVC Class E KristenAndSara –22.8 –43.1 videoSRC14 –48.6 –44.5
FourPeople –28.8 –42.4 videoSRC15 –30.2 –28.7
Johnny –20.5 –43.3 videoSRC16 –11.8 –19.4
HEVC Class RGB EBULupoCandlelight –19.4 –28.3 videoSRC17 –29.1 –29.6
OldTownCross –33.5 –35.0 videoSRC18 12.5 16.4
ParkScene 0.3 –16.2 videoSRC19 –30.0 –6.8
DucksAndLegs –18.0 –7.9 videoSRC20 50.2 68.2
Kimono1 –29.6 –29.6 videoSRC21 –25.7 –17.1
EBURainFruits –6.0 –35.8 videoSRC22 –26.3 –24.0
UVG HoneyBee –21.7 –47.2 videoSRC23 –37.8 –31.6
ReadySteadyGo –31.2 –18.1 videoSRC24 –33.4 –30.8
Bosphorus –1.0 –29.5 videoSRC25 –21.7 –23.2
YachtRide –26.8 –26.5 videoSRC26 –26.6 –23.7
Jockey –3.0 90.9 videoSRC27 –11.4 –23.7
ShakeNDry –21.7 –24.3 videoSRC28 –20.5 –31.3
Beauty –61.3 –52.4 videoSRC29 0.2 –11.7
- - - - videoSRC30 –13.3 –35.8
@@ V-C5 Influence of Larger GOP Size
To evaluate the capability of our DCVC-B in supporting a larger GOP size, we conduct an ablation study by setting its intra period and GOP size to 64. It is important to note that we did not fine-tune our model using longer training sequences. For comparison, we also set the intra period of DCVC-DC to 64. In Fig. 16, we compare the frame quality and bitrate of our DCVC-B scheme and those of DCVC-DC. The comparison results verify that our DCVC-B scheme can effectively support a larger GOP size and mitigate the error propagation problem.
@@ VI Limitations and Future Work
As reported in Table II, our proposed DCVC-B exhibits different compression performance across various testing datasets. In this section, we aim to identify the limitations of DCVC-B by analyzing the characteristics of each dataset. We compare the BD-rate values of DCVC-DC and DCVC-B against the anchor HM-RA-GOP16 for each video sequence in Table VII. The comparison results reveal that DCVC-B is inferior to DCVC-DC when the testing videos have larger motion, as it struggles to obtain accurate bi-directional temporal predictions. Conversely, DCVC-B is superior to DCVC-DC when the testing videos have smaller motion, as it can leverage more useful temporal information from accurate bi-directional temporal predictions. For example, as shown in Fig. 17, when the motion between bi-directional reference frames is large, DCVC-B can only predict blurred bi-directional temporal contexts for sequences such as the $BasketballDrive\_1920\times 1080\_50$ sequence in the HEVC Class B dataset and the $RaceHorses\_832\times 480\_30$ sequence in the HEVC Class C dataset. In this case, it is difficult for DCVC-B to leverage useful temporal information from bi-directional reference frames, leading to lower compression performance. In contrast, when the motion between reference frames is small, DCVC-B can predict accurate bi-directional temporal contexts, such as the $BQSquare\_416\times 240\_60$ sequence in the HEVC Class D dataset and the $FourPeople\_1280\times 720\_60$ sequence in the HEVC Class E dataset. In this case, with more useful temporal information from bi-directional temporal contexts, DCVC-B can effectively reduce the conditional entropy, leading to higher compression performance.
@@ VII Conclusion
In this paper, we propose a bi-directional deep contextual video compression scheme tailored for B-frames, termed DCVC-B. We focus on improving the compression performance for the B-frame coding from three aspects. Firstly, we propose a bi-directional motion difference context propagation method to reduce the motion coding costs. Secondly, we propose a bi-directional contextual compression model and a bi-directional temporal entropy model to make better use of multi-scale temporal contexts. Thirdly, we propose a hierarchical quality structure-based training strategy to achieve a better bit allocation within a large GOP. Experimental results demonstrate that, in terms of PSNR, our DCVC-B scheme significantly outperforms the reference software of H.265/HEVC under the random access configuration. On some testing datasets, DCVC-B even surpasses the reference software of H.266/VVC. Additionally, we analyze the unique challenges of deep B-frame coding and identify the limitations of our scheme. To address these limitations, we propose important future work that could further enhance the compression performance of deep B-frame coding, thereby beneficial for the community.

### v4

@@ II-A Deep Video Compression for P-Frame
For the second aspect, Li et al. [8] proposed a deep contextual video compression scheme, termed DCVC, which replaces the residual coding paradigm with a conditional coding paradigm. Instead of relying on the subtraction operation to reduce temporal redundancy, DCVC regards the temporal prediction as a condition so that the codec can learn how to make full use of the temporal prediction automatically. To further utilize temporal predictions, Sheng et al. [11] proposed DCVC-TCM, which not only uses temporal predictions at the entrance of the contextual encoder but also feeds the multi-scale temporal predictions into its intermediate locations. In addition, they also proposed to learn a temporal prior from the multi-scale temporal predictions to utilize temporal predictions in the entropy model. Based on DCVC-TCM, Li [9] proposed DCVC-HEM, which introduced a latent prior into the temporal entropy model to further utilize temporal correlation.
@@ IV-E Hierarchical Quality Structure
To achieve a hierarchical quality structure for B-frame coding, traditional video codecs commonly assign different QPs (quantization parameters) to video frames of different temporal layers [55]. Recent deep P-frame coding schemes [6, 10] periodically increase the quality of video frames by assigning a quality coefficient that varies periodically over the frame index to the Lagrangian multiplier $\lambda$ in the rate-distortion (R-D) loss function. Inspired by traditional video codecs and deep P-frame coding schemes, we propose to assign different quality coefficients $w_{t}$ for the B-frames of different temporal layers. As shown in Fig. 2, we set $w_{t}$ of the B-frames in Layer 1, 2, 3, 4, and 5 to [1.4, 1.4, 0.7, 0.5, 0.5], respectively. When calculating the loss functions to train our model, we multiply $w_{t}$ on the $\lambda$. In addition, to make our model adapt to different qualities better, a quality adaptor implemented by one $1\times 1$ convolutional layer is inserted before the feature extractor module of the temporal context mining. For B-frames of different temporal layers, different quality adaptors are selected. Note that the quality adaptors are included in the training from the beginning.
We combine the hierarchical quality structure-based training strategy with the step-by-step training strategy. The detailed training strategies are listed in Table. I. For different numbers of training frames, the training reference structures are illustrated in Fig. 8. The loss functions for different steps are defined as follows.
@@
Unlike existing deep P-frame coding schemes [11, 9, 10, 12, 6], we do not calculate the average loss for multi-frame joint training. When using $L_{t}^{meD}$ or $L_{t}^{meRD}$ as the loss function, we only train the motion-related modules (Inter). When using $L_{t}^{recD}$ or $L_{t}^{recRD}$ as the loss function, we only train the context-related modules (Recon). When using $L_{t}^{all}$ as the loss function, we jointly train all the modules (All).
@@ V-A4 Test Configurations
For neural video coding models, we compare with existing state-of-the-art deep P-frame coding schemes, including DCVC-HEM [9], DCVC-SDD [6], DCVC-DC [10], and DCVC-FM [12]. We also compare with existing deep B-frame coding scheme B-CANF [14]. For our scheme, we follow the existing deep P-frame coding schemes [9, 6, 10, 12] and compress 96 frames of one video to make a fair comparison. For B-CANF, we use its default setting and compress 97 frames but calculate BD-rate using the compression results of the first 96 frames. The intra period and GOP size of our DCVC-B scheme and existing deep coding schemes are all set to 32. When compressing 96 frames, the structure of the first two GOPs of our codec is shown in Fig. 2 and the structure of the last GOP is shown in Fig. 9, which is the same as the VTM-RA-GOP32 except that it has only 2 reference frames (VTM-RA-GOP32 has 4 reference frames). We use the same I-frame codec as DCVC-DC.
@@ V-B1 Comparison Results
We illustrate the RD-curves in terms of PSNR and MS-SSIM for different testing datasets in Fig. 10 and Fig. 11. We also report the corresponding BD-rate values in Table II and Table III. The anchor is HM-RA-GOP16. Negative values indicate bitrate saving compared with HM while positive values indicate bitrate increasing. The comparison results show that in terms of PSNR, our proposed DCVC-B scheme obtains an average –26.6% BD-rate reduction against the anchor HM-RA-GOP16, which even outperforms the SOTA deep P-frame coding schemes DCVC-DC (–24.4%) and DCVC-FM (–23.4%). For the HEVC Class D and Class E datasets, we even outperform VTM-RA-GOP32. Compared with the state-of-the-art deep B-frame coding scheme B-CANF [14], our DCVC-B scheme achieves a large compression performance improvement. In terms of MS-SSIM, DCVC-B obtains an average –49.9% BD-rate reduction against HM-RA-GOP16 and even outperforms VTM-RA-GOP32 setting (–34.3%) over all the datasets. The subjective comparison results illustrated in Fig. 12 show that the reconstructed videos of our scheme can retain more details.
@@ V-C4 Effectiveness of Hierarchical Quality Structure
Figure 16: Frame quality and bitrate comparison between our proposed DCVC-B, DCVC-DC, and DCVC-FM when the GOP size is set to 64.
@@
Datasets Sequences DCVC-DC DCVC-FM Ours Datasets Sequences DCVC-DC DCVC-FM Ours
HEVC B Cactus –31.0 –28.9 –37.2 MCL-JCV videoSRC01 –17.4 –28.4 –33.0
ParkScene –9.9 –6.3 –30.8 videoSRC02 –16.0 –12.0 0.4
Kimono1 –29.2 –27.8 –29.7 videoSRC03 –23.4 –20.6 –37.4
BQTerrace –12.2 –18.4 –20.5 videoSRC04 –19.5 –16.2 –13.4
BasketballDrive –26.3 –26.9 –11.8 videoSRC05 –30.0 –21.4 –11.3
HEVC C BQMall –20.3 –22.7 –14.8 videoSRC06 –28.8 –46.9 56.1
BasketballDrill –25.1 –24.0 –26.9 videoSRC07 –29.9 –20.1 –29.9
RaceHorses –22.0 –20.6 –0.8 videoSRC08 –31.7 –27.1 –20.8
PartyScene –15.4 –16.7 –23.7 videoSRC09 –26.6 –25.9 –40.1
HEVC D BasketballPass –36.4 –36.8 –38.4 videoSRC10 –13.6 4.7 4.6
BQSquare –22.0 –25.1 –42.4 videoSRC11 –35.5 –25.4 –7.6
BlowingBubbles –24.5 –22.7 –37.3 videoSRC12 –1.3 –2.5 –31.3
RaceHorses –42.4 –37.8 –34.9 videoSRC13 19.6 3.8 –18.0
HEVC E KristenAndSara –22.8 –27.5 –43.1 videoSRC14 –48.6 –48.2 –44.5
FourPeople –28.8 –34.1 –42.4 videoSRC15 –30.2 –28.4 –28.7
Johnny –20.5 –29.2 –43.3 videoSRC16 –11.8 –18.9 –19.4
HEVC RGB EBULupoCandlelight –19.4 –21.7 –28.3 videoSRC17 –29.1 –27.9 –29.6
OldTownCross –33.5 –26.5 –35.0 videoSRC18 12.5 19.8 16.4
ParkScene 0.3 15.0 –16.2 videoSRC19 –30.0 –25.9 –6.8
DucksAndLegs –18.0 5.0 –7.9 videoSRC20 50.2 73.0 68.2
Kimono1 –29.6 –27.8 –29.6 videoSRC21 –25.7 –22.1 –17.1
EBURainFruits –6.0 –11.3 –35.8 videoSRC22 –26.3 –24.7 –24.0
UVG HoneyBee –21.7 –16.8 –47.2 videoSRC23 –37.8 –40.0 –31.6
ReadySteadyGo –31.2 –27.2 –18.1 videoSRC24 –33.4 –25.9 –30.8
Bosphorus –1.0 –7.9 –29.5 videoSRC25 –21.7 –17.7 –23.2
YachtRide –26.8 –29.6 –26.5 videoSRC26 –26.6 –19.9 –23.7
Jockey –3.0 23.2 90.9 videoSRC27 –11.4 –13.4 –23.7
ShakeNDry –21.7 –23.5 –24.3 videoSRC28 –20.5 –24.9 –31.3
Beauty –61.3 –65.4 –52.4 videoSRC29 0.2 –23.1 –11.7
- - - - - videoSRC30 –13.3 –11.5 –35.8
@@ V-C5 Influence of Larger GOP Size
To evaluate the capability of our DCVC-B to support a larger GOP size, we conduct an ablation study by setting its intra period and GOP size to 64. It is important to note that we did not fine-tune our model using longer training sequences. For comparison, we also set the intra period of DCVC-DC and DCVC-FM to 64. In Fig. 16, we compare the frame quality and bitrate of different codecs. Observing the comparison results, we find that DCVC-FM has less error propagation than DCVC-DC because it uses 32 frames for joint training and uses the feature re-fresh method. Without the need for multi-frame joint training that consumes a lot of GPU memory, our DCVC-B can rely on the bi-directional hierarchical quality structure to reduce error propagation.
@@ VI Limitations and Future Work
As reported in Table II, our proposed DCVC-B exhibits different compression performance across various testing datasets. In this section, we aim to identify the limitations of DCVC-B by analyzing the characteristics of each dataset. We compare the BD-rate values of DCVC-DC, DCVC-FM, and DCVC-B against the anchor HM-RA-GOP16 for each video sequence in Table VII. The comparison results reveal that DCVC-B is inferior to DCVC-DC and DCVC-FM when the testing videos have larger motion, as it struggles to obtain accurate bi-directional temporal predictions. Conversely, DCVC-B is superior to DCVC-DC and DCVC-FM when the testing videos have smaller motion, as it can leverage more useful temporal information from accurate bi-directional temporal predictions. For example, as shown in Fig. 17, when the motion between bi-directional reference frames is large, DCVC-B can only predict blurred bi-directional temporal contexts for sequences such as the $BasketballDrive\_1920\times 1080\_50$ sequence in the HEVC Class B dataset and the $RaceHorses\_832\times 480\_30$ sequence in the HEVC Class C dataset. In this case, it is difficult for DCVC-B to leverage useful temporal information from bi-directional reference frames, leading to lower compression performance. In contrast, when the motion between reference frames is small, DCVC-B can predict accurate bi-directional temporal contexts, such as the $BQSquare\_416\times 240\_60$ sequence in the HEVC Class D dataset and the $FourPeople\_1280\times 720\_60$ sequence in the HEVC Class E dataset. In this case, with more useful temporal information from bi-directional temporal contexts, DCVC-B can effectively reduce the conditional entropy, leading to higher compression performance.
@@ VII Conclusion
In this paper, we propose a bi-directional deep contextual video compression scheme tailored for B-frames, termed DCVC-B. We focus on improving the compression performance for the B-frame coding from three aspects. Firstly, we propose a bi-directional motion difference context propagation method to reduce the motion coding costs. Secondly, we propose a bi-directional contextual compression model and a bi-directional temporal entropy model to make better use of multi-scale temporal contexts. Thirdly, we propose a hierarchical quality structure-based training strategy to achieve a better bit allocation within a large GOP. Experimental results demonstrate that, in terms of PSNR, our DCVC-B scheme significantly outperforms the reference software of H.265/HEVC under the random access configuration. On some testing datasets, DCVC-B even surpasses the reference software of H.266/VVC. Additionally, we analyze the unique challenges of deep B-frame coding and identify the limitations of our scheme. To address these limitations, we propose important future work that could further enhance the compression performance of deep B-frame coding, thereby beneficial for the community. With the gradual improvement of deep P-frame codecs in recent years, we believe that deep B-frame codecs will be better and better.